# From multitype branching Brownian motions to branching Markov additive processes

### v1

@@ 5.3 The spine decomposition with respect to the truncated derivative martingale
$\mathrm{d}\widetilde{\mathbf{P}}^{\uparrow}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}$
$=$ $\zeta_{t}^{(2)\uparrow}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\zeta_{t}^{(1)}\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\zeta_{t}^{(3)}\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})\prod_{v\prec\xi_{t}}\bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\bigg]$
$=$ $\mathrm{d}\mathbb{P}_{x,i}^{\uparrow}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}\prod_{v\prec\xi_{t}}\bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\bigg].$

### v2

@@ 5.3 The spine decomposition with respect to the truncated derivative martingale
$\mathrm{d}\widetilde{\mathbf{P}}^{\uparrow}_{x,i}(\tau,M,\xi)|_{\widetilde{\mathcal{F}}_{t}}=$ $\,\zeta_{t}^{(2)\uparrow}\mathrm{d}\mathbb{P}_{x,i}\left((X_{\xi},J_{\xi})_{t}\right)\zeta_{t}^{(1)}\mathrm{d}L^{\beta(J_{\xi})}(\mathbf{n}_{t})\zeta_{t}^{(3)}\prod_{v\prec\xi_{t}}\mu_{J_{\xi}(d_{v})}(A_{v})$
$\,\prod_{v\prec\xi_{t}}\bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\bigg]$
$=$ $\,\mathrm{d}\mathbb{P}_{x,i}^{\uparrow}\left((X_{\xi},J_{\xi})_{t}\right)\mathrm{d}L^{(m\beta)(J_{\xi})}(\mathbf{n}_{t})\prod_{v\prec\xi_{t}}\frac{\mu_{J_{\xi}(d_{v})}(A_{v})A_{v}}{m(J_{\xi}(d_{v}))}$
$\,\prod_{v\prec\xi_{t}}\bigg[\frac{1}{A_{v}}\prod_{j:vj\in O_{v}}\mathrm{d}\mathbf{P}_{X_{\xi}(d_{v}),J_{\xi}(d_{v})}((\tau,M)_{t-d_{v}}^{v,j})\bigg].$